OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = -3(x + 8)2 + 5?

OpenStudy (jamesj):

What's the axis of symmetry of y = x^2 ? It's the y-axis, or the line x = 0 What's the axis of symmetry of y = ax^2? The same, the line x = 0. The axis of symmetry of y = ax^2 + c? The same, x = 0 Axis of symmetry of y =(x-1)^2 ? x = 1 Of y = a(x-1)^2 ? x = 1 Of y = a(x-1)^2 + b? x = 1

OpenStudy (jamesj):

What then is the axis of symmetry of y = (x + 8)^2 ?

OpenStudy (jamesj):

hint: (x+8)^2 = ( x - (-8) )^2

OpenStudy (jamesj):

any ideas?